mathslover (mathslover):

Starting Integral Calculus Indefinite Integral (Algebra of Indefinite Integral) If \(\int f(x) dx = F(x) + C\) , then \(\int f(ax+b) dx = \cfrac{1}{a} F(ax+b) + C \) Is there any way to prove it?

OpenStudy (anonymous):

yep there is :) for example :- differentiate both side w.r.t X

OpenStudy (anonymous):

(1/a F(ax+b) ) ' = 1/a (ax+b )' F(ax+b)' =1/a *a * F(ax+b)' = F(ax+b)'

mathslover (mathslover):

Perfect! Got it. Thanks!

OpenStudy (anonymous):

YW :)